OpenStudy (cenaida):

Evaluate 56 divided by 53. A. 125 B. 0 C. 5 D. 25

OpenStudy (cenaida):

5^6 divided by 5^3

OpenStudy (anonymous):

5^3=125

OpenStudy (anonymous):

\[\frac{5^6}{ 5^3 } = 5^{6-3}=5^3= 5 \times 5 \times 5 \]

hartnn (hartnn):

\(\huge \dfrac{x^m}{x^n}=x^{m-n}\)

OpenStudy (anonymous):

@hartnn How do you gets such a big font ?

OpenStudy (cenaida):

good question haha

hartnn (hartnn):

use `\huge` while writing in latex `\huge \dfrac{x^m}{x^n}=x^{m-n}`

OpenStudy (anonymous):

@hartnn Thank you